# Galois Action and Localization in Number Fields

### v2

@@ Abstract.
For a Galois number field $K$, the Galois group $\text{Gal}(K/\mathbb{Q})$ acts on the class group $Cl_{K}$ in a very natural way: $\sigma\cdot[I]=[\sigma(I)]$ for any $\sigma\in\text{Gal}(K/\mathbb{Q})$, $I]\in Cl_{K}$. In this paper, we will explore how the unique properties of this group action work together to elucidate the relationship between these two groups. While previous work on this problem has focused on representation theory, we take a direct approach to some classical and new problems. The paper concludes with an exploration of the class groups of overrings of the ring of integers $\mathcal{O}_{K}$. These turn out to be powerful tools for understanding $Cl_{K}$ and overrings of $\mathcal{O}_{K}$.
@@ 1. Introduction and Notation
Throughout this paper, we will see how properties 1-4 can be used to place strong restrictions on the structure of $Cl_{K}$ given $G$ and vice versa. We will place a particular emphasis on techniques developed and some applications to factorization theory.
@@
The previous work done on this subject has been primarily through the lens of representation theory—considering $Cl_{K}$ as a $G$-module. This approach was utilized by Fröhlich in 1952 ([9]) and subsequently many others. Cornell and Rosen used this approach in [6] to give results on the structure of $Cl_{K}$ when $h_{K}$ is known. Others such as Lemmermeyer ([12]) and Iwasawa ([11]) used the $G$-module structure of $Cl_{K}$ along with assumptions on intermediate fields to give various results on the $p$-rank of the class group. For a thorough review of the previous work done in this area, see [14].
@@
In the following section, we will take a closer look at the conditions of Definition 1.1 and their implications for norm-like actions in general. In Section 3, we explore how conditions 1-2 together with 4 place restrictions of the structure of $Cl_{K}$ for a Galois number field of given degree. In Section 4, we focus on applications of condition 3 to the inverse class group problem. In particular, which number fields $K$ can have a given abelian group as their class group. In Section 5, the authors determine the structure of the class group for localizations of the form $\mathcal{O}_{K}[\frac{1}{x}]$ which ultimately strengthens many of the previously developed results and leads to some interesting observations about rings of integers. Finally, Section 6 concludes with a more general exploration of overrings of Dedekind domains, culminating with a characterization of those overrings whose class groups admit a norm-like group action.
2. Norm-Like Group Actions
In this section, we will consider norm-like group actions directly. The results and observation made here will be used throughout the remainder of the paper. First, let us recall the definitions of a norm-like action.
Definition 1.1.
Let $G$ and $A$ be groups with $A$ abelian, and let
$\alpha:G\times A\longrightarrow A,\quad(g,a)\mapsto g\cdot a$
be a map. If $\alpha$ satisfies the following properties:
  1. (1)
    $g_{1}\cdot(g_{2}\cdot a)=(g_{1}g_{2})\cdot a$ for all $g_{1},g_{2}\in G$, $a\in A$,
  2. (2)
    $e_{G}\cdot a=a$ for all $a\in A$,
  3. (3)
  4. (4)
    $\prod_{g\in G}(g\cdot a)=e_{A}$ for all $a\in A$,
then we say that $\alpha$ is a norm-like action.
Conditions $1$ and $2$ are those of a typical group action. The definition diverges with condition $3$ which ensures the action of each element $g$ on $A$ behaves like a homomorphism. In fact, each element of $G$ can be identified with an endomorphism of $A$ via the map $g\mapsto\phi$ where $\phi(a)=g\cdot a$. If $A$ is finite, in the case where $g\cdot a=e_{A}\implies a=e_{A}$, this map defines a group homomorphism from $G$ to $\text{Aut}(A)$, the kernel of which is the set $\{g\in G\,|\,g\cdot a=a\,\forall\,a\in A\}$. Note, this is precisely the kernel of the group action. The implications of this are closely considered in Section 4. We also note that condition $3$ guarantees $g\cdot e_{A}=e_{A}$ for all $g\in G$. Thus, a norm-like group action will be transitive if and only if $A$ is trivial.
Let us consider a group action of $G$ on $A$ satisfying conditions 131-3. First, it suffices to check condition $4$ for the generators of $A$. Now, consider the subgroup $H:=\{\prod_{g\in G}g\cdot a\,|\,a\in A\}\unlhd A$. For any $x\in H$, we have $x=\prod_{g\in G}g\cdot a^{\prime}$ for some $a^{\prime}\in A$. Thus, for any $g^{\prime}\in G$, $g^{\prime}\cdot x=g^{\prime}\cdot\left(\prod_{g\in G}g\cdot a^{\prime}\right)=\prod_{g\in G}(g^{\prime}*g)\cdot a^{\prime}=\prod_{g\in G}g\cdot a^{\prime}=x$. Hence, $H$ is a subgroup of the elements of $A$ with trivial orbit. Therefore, if $e_{A}$ is the only element with a trivial orbit, condition $4$ follows from 131-3.
Now, let $A$ be an arbitrary abelian group, and $G=\{id,-id\}\leqslant\text{Aut}(A)$ where $id$ is the identity, and id(a)=a-id(a)=-a for all $a\in A$ which is a automorphism just in case $A$ is abelian. It is easy to verify that the natural action $g\cdot a=g(a)$ is norm-like. This case has been of particular interest over the last three years as the study of weighted zero sum sequences has become popular in connection with normset factorization (see [1] and [8]). In particular, $\{id,-id\}$ is the subgroup induced by the action of $\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$ when $K$ is quadratic. As we have seen, $|G|=2$ places no restriction on $A$. As we will see throughout this paper, this is a very unique case. Even the $|G|=3$ case is much more restrictive and the relevant arithmetic far more complicated.
Example 2.1.
Cubic Example: $G\cong\mathbb{Z}/3\mathbb{Z}$ acting on $A\cong\mathbb{Z}/7\mathbb{Z}$ Also, any of the subgroups of $SL_{2}(\mathbb{Z}/2\mathbb{Z})$ (order $6$) acting on $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ seem to work. There might be something here. Probably want a non-example in the cubic case as well.Is this unique in the sense that the action of all subgroups of $\text{Aut}(A)$ work via the natural group action.
@@ 3. Direct Consequences of the Group Action
Let $p<23$ be an odd prime and $a\in\mathbb{Z}$ not divisible by the $p^{th}$ power of any prime. If $K$ is the splitting field of $x^{p}-a$, $\mathcal{O}_{K}$ is an HFD if and only if it is $\mathcal{O}_{K}$ is a UFD.
@@
We first note that, as the splitting field of $x^{p}-a$, $K$ is Galois over $\mathbb{Q}$. Furthermore, we have $K\cong\mathbb{Q}(\omega,\alpha)$ where $\omega$ is a primitive $p^{th}$ root of unity and $\alpha$ a root of $x^{p}-a$. This gives us the following lattice.
@@
Recall that the action of $G=\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$ induces a homomorphism from $G$ to $\text{Aut}(Cl_{K})$ given by $\sigma\overset{\psi}{\mapsto}\phi$ where $\phi([I])=\sigma\cdot[I]=[\sigma(I)]$. In this section, we will see how this map can be used to further restrict the structure of $Cl_{K}$ for a Galois number field $K$. Conversely, we will use this tool to help answer questions related to the inverse class group problem. This terminology is used in the spirit of the inverse Galois problem. In particular, the inverse class group problem asks if an arbitrary abelian group can be realized as the class group of a ring of integers. This question was famously answered in the positive for the more general class of Dedekind domains by Claborn in [5]. In the case of imaginary quadratics, the answer is known to be no. Certain groups can be shown by brute force not to appear, the smallest among them being $(\mathbb{Z}/3\mathbb{Z})^{3}$. The question remains open for rings of integers in general. In this section, for a given finite abelian group $A$, we will explore which rings of integers may admit $Cl_{K}\cong A$. We begin with a direct proof of another result from [6] which we will return to with some new found tools in section 4.
@@
5. Localizations of Rings of Integers
In this section, we will use localization to strengthen some of the results developed previously. First, recall that any overring—in particular, any localization—of a Dedekind domain remains Dedekind. Intuitively, as a localization of $R$ is attained by turning a set of elements $S\subseteq R$ into units, we expect the class group of $R_{S}$ to be smaller than that of $R$ as those non-principal ideals $I\subseteq R$ for which $I\cap S\neq\emptyset$ will become principal. As we will demonstrate, “small” localizations will tend to lead to small changes in the class group. In the context of rings of algebraic integers, this will be useful in further restricting the possible structure of $Cl_{K}$.
@@ 5. Localizations of Rings of Integers
Let $D$ be a Dedekind domain and $x\in D$ a nonzero nonunit with
$(x)=\mathfrak{p}_{1}^{n_{1}}\mathfrak{p}_{2}^{n_{2}}\dotsm\mathfrak{p}_{k}^{n_{k}}$
Then $Cl(D[\frac{1}{x}])\cong Cl(D)/\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$>.
Note, $D[\frac{1}{x}]$ is the localization of $D$ at the multiplicative set $\{x^{n},\,n\in\mathbb{N}\}$ which is the complement of the set-theoretic union of all prime ideals not containing $x$.
@@
Let $\phi:Cl(D)\rightarrow Cl(D[\frac{1}{x}])$ be the canonical map $I]\overset{\phi}{\mapsto}\left[ID[\frac{1}{x}]\right]$. First we will show that $\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$> is contained in $\text{ker}(\phi)$. It suffices to show that $\mathfrak{p}_{i}$ becomes principal in $D[\frac{1}{x}]$ for any $1\leqslant i\leqslant k$. This is clearly the case as $x\in\mathfrak{p}_{i}$, so $\mathfrak{p}_{i}D[\frac{1}{x}]=D[\frac{1}{x}]$.
Now, let $I]\in\text{ker}(\phi)$. The case when $I]=\text{Prin}(D)$ is trivial, so we will assume $I$ is a non-principal, integral ideal. As $D$ is Dedekind, this implies $I$ is $2$-generated, so let us write $I=(r_{1},r_{2})D$. Now, $I]\in\text{ker}(\phi)$ implies $ID[\frac{1}{x}]$ is principal, so we write $ID[\frac{1}{x}]=(r_{1},r_{2})D[\frac{1}{x}]=rD[\frac{1}{x}]$. Note, we may assume without loss of generality that $r\in D$ because for any $\frac{a}{x^{n}}\in D[\frac{1}{x}]$, $\frac{a}{x^{n}}D[\frac{1}{x}]=aD[\frac{1}{x}]$. Now, the equality implies there exist $\frac{s_{1}}{x^{n_{1}}},\frac{s_{2}}{x^{n_{2}}}\in D[\frac{1}{x}]$ such that $r\frac{s_{1}}{x^{n_{1}}}=r_{1}$ and $r\frac{s_{2}}{x^{n_{2}}}=r_{2}$. Also, we must have $\frac{t_{1}}{x^{m_{1}}},\frac{t_{2}}{x^{m_{2}}}\in D[\frac{1}{x}]$ such that $r=r_{1}\frac{t_{1}}{x^{m_{1}}}+r_{2}\frac{t_{2}}{x^{m_{2}}}$. Without loss of generality, assume $n_{1}\geqslant n_{2}$ and $m_{1}\geqslant m_{2}$. Then, as $rs_{1}=r_{1}x^{n_{1}}$ and $rs_{2}=r_{2}x^{n_{2}}$, we see that $r$ divides $x^{n_{1}}r_{1}$ and $x^{n_{1}}r_{2}$ in $D$. Thus, $J=(\frac{x^{n_{1}}r_{1}}{r},\frac{x^{n_{1}}r_{2}}{r})D$ is an integral ideal of $D$, and $I\sim J$ as they differ by the principal ideal $(\frac{x^{n_{1}}}{r})D$. Now, from $r=r_{1}\frac{t_{1}}{x^{m_{1}}}+r_{2}\frac{t_{2}}{x^{m_{2}}}$ we get $1=t_{1}\frac{r_{1}}{rx^{m_{1}}}+t_{2}\frac{r_{1}}{rx_{m_{2}}}$, and multiplying by $x^{n_{1}+m_{1}}$, we get $x^{n_{1}+m_{1}}=t_{1}\frac{x^{n_{1}}r_{1}}{r}+t_{2}x^{m_{1}-m_{2}}\frac{x^{n}_{1}r_{2}}{r}\in J$.
The $\mathfrak{p}_{i}$ which divide $(x)$ are precisely those prime ideals which contain $x$, so we must have $J=\mathfrak{p}_{1}^{a_{1}}\mathfrak{p}_{2}^{a_{2}}\dotsm\mathfrak{p}_{k}^{a_{k}}$ with the possibility that $a_{i}=0$. Therefore, $I]=[J]=[\mathfrak{p}_{1}]^{a_{1}}[\mathfrak{p}_{2}]^{a_{2}}\dotsm[\mathfrak{p}_{k}]^{a_{k}}\in\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$>, so the result follows from the first isomorphism theorem.
@@
Notably, given $Cl(D)$ is torsion, Theorem 5.1 implies that any homomorphic image of $Cl(D)$ can be realized as the class group of such a localization of $D$. Claborn gives a similar, slightly weaker result in [4] where, given a Dedekind domain $A$ with arbitrary class group $G$ and subgroup $H$, he constructs a Dedekind polynomial ring over $A$ with class group isomorphic to $G/H$. As he answered the inverse class group problem for Dedekind domains, Theorem 5.1 brings us a step closer for overrings of $\mathcal{O}_{K}$. In essence, $\mathcal{O}_{K}[\frac{1}{x}]$ is very nearly a ring of integers itself. In fact, it is the precise analog if the base ring $\mathbb{Z}$ were replaced by $\mathbb{Z}[\frac{1}{x}]$ for some nonzero $x$ in the number field $K$. Note also that $Cl(\mathcal{O}_{K}[\frac{1}{x}])$ is finite and retains the useful property that each class contains a prime ideal. This follows directly from the correspondence between prime ideals of $\mathcal{O}_{K}[\frac{1}{x}]$ and prime ideals of $\mathcal{O}_{K}$ disjoint from $\{x^{n},\,n\in\mathbb{N}\}$. This allows for the application of many theorems on rings of integers to be applied to $\mathcal{O}_{K}[\frac{1}{x}]$. The following gives one such example.
@@
The remainder of this section will be dedicated to leveraging this fact to extend the techniques and theorems developed in sections 2 and 3. Our hope is that the Galois group of $K$ will act on the class group of $\mathcal{O}_{K}[\frac{1}{x}]$ in the same manner as $\text{Cl}_{K}$ which would allow us to place the same restrictions on the homomorphic images of the class group—further constraining the structure of $Cl_{K}$. Unfortunately, the action $\sigma\cdot[\mathfrak{P}]=[\sigma(\mathfrak{P})]$ for $\sigma\in G$ and $\mathfrak{P}]\in Cl(\mathcal{O}_{K}[\frac{1}{x}])$ will not be well-defined in general as $\sigma(\frac{1}{x})=\sigma(x)^{-1}$ need not be in $\mathcal{O}_{K}[\frac{1}{x}]$ in general. However, in the case that $x\in\mathbb{Z}$, we avoid this issue as we will have $\sigma(\frac{1}{x})=\frac{1}{x}$ for all $\sigma\in G$.
@@
Let $K$ be a Galois number field and $A$ the subgroup of $\text{Aut}(Cl_{K})$ induced by the action of of $\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$. Suppose that $x\in\mathcal{O}_{K}$ is a nonzero nonunit and
$(x)=\mathfrak{p}_{1}^{n_{1}}\mathfrak{p}_{2}^{n_{2}}\dotsm\mathfrak{p}_{k}^{n_{k}}.$
Then $Cl(\mathcal{O}_{K}[\frac{1}{N(x)}])\cong Cl_{K}/A\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$> where $N(x)$ is the norm of $x$, and $A\left<X\right$> denotes the subgroup generated by the $A$-automorphic images of the elements in $X$.
@@
Recall that the ideal norm agrees with the element norm for principal ideals. Thus, as $K$ is Galois, we have $(N(x))=N((x))=\prod_{\sigma\in G}\sigma((x))=\prod_{\sigma\in G}\sigma(\mathfrak{p}_{1})^{n_{1}}\dotsm\sigma(\mathfrak{p}_{k})^{n_{k}}$. The theorem then follows by applying Theorem 5.1 to $N(x)$. ∎
Now, the group action of $G$ on the class group of $\mathcal{O}_{K}[\frac{1}{N(x)}]$ will be well-defined. More than this, it will be a Galois action.
@@
For a Galois number field $K$ and $x\in\mathcal{O}_{K}$ a nonzero nonunit, the group action $\sigma\cdot[\mathfrak{P}]=[\sigma(\mathfrak{P})]$ for $\sigma\in\text{Gal}(F/\mathbb{Q}),[\mathfrak{P}]\in Cl\left(\mathcal{O}_{K}[\frac{1}{N(x)}]\right)$ is a norm-like action.

### v4

@@ Abstract.
For a Galois number field $K$, the Galois group $\text{Gal}(K/\mathbb{Q})$ acts on the class group $Cl_{K}$ in a very natural way: $\sigma\cdot[I]=[\sigma(I)]$ for any $\sigma\in\text{Gal}(K/\mathbb{Q})$, $I]\in Cl_{K}$. In this paper, we will explore how the unique properties of this group action work together to elucidate the relationship between these two groups—developing and expanding upon some known results from a new perspective. To this end, we explore the class groups of localizations of the ring of integers $\mathcal{O}_{K}$. These turn out to be powerful tools for understanding $Cl_{K}$ and overrings of $\mathcal{O}_{K}$. The paper concludes with some interesting observations about normset arithmetic—a topic intimately related to this action.
@@ 1. Introduction and Notation
Throughout this paper, we will see how properties 1-4 can be used to place strong restrictions on the structure of $Cl_{K}$ given $G$ and vice versa. We will place a particular emphasis on techniques developed and some applications to factorization theory. While the class group is inherently connected to the arithmetic of rings of integers, this particular action is intimately related to the factorization properties of the normset of a ring of integers. This connection was first studied by the first author in [11]. The recent work of Geroldinger, Halter-Koch, and Zhong in [15] has inspired renewed interest in this action as it relates to normset factorization—particularly in the quadratic case.
@@
Throughout this paper, we will present new proofs of some existing theorems as well as extend various results on the structure of $Cl_{K}$ using new techniques. Previous work done on this subject has been primarily through the lens of representation theory—considering $Cl_{K}$ as a $G$-module. This approach was utilized by Fröhlich in 1952 ([14]) and subsequently many others. Cornell and Rosen used this approach in [10] to give results on the structure of $Cl_{K}$ when $h_{K}$ is known. Others such as Lemmermeyer ([21]) and Iwasawa ([19]) used the $G$-module structure of $Cl_{K}$ along with assumptions on intermediate fields to give various results on the $p$-rank of the class group. In [20], Komatsu and Nakano generalize this work, using the Galois-module structure of the class group and some group theoretic considerations, to give further restrictions on the $p$-rank of $Cl_{K}$—once again using assumptions the class group of intermediate fields. We will develop some of the tools used in this paper from a strictly group-theoretic point of view. However, we diverge from these methods by considering the class groups of overring of $\mathcal{O}_{K}$ rather than that of intermediate rings. For a thorough review of the previous work done in this area, see [23].
@@
In the following section, we explore how conditions 1 and 2, the typical conditions of a group action, together with 4 place restrictions of the structure of $Cl_{K}$ for a Galois number field $K$ of given degree. In Section 3, we focus on applications of condition 3 to the inverse class group problem. In particular, which number fields $K$ can have a given abelian group as their class group. Section 4 considers the structure of the class group of localizations of the form $\mathcal{O}_{K}[\frac{1}{\alpha}]$. This ultimately strengthens many of the previously developed results by allowing us to restrict to Sylow $p$-subgroups of the class group. In Section 5, the authors consider the class groups of overrings of Dedekind domains broadly. The paper culminates with a characterization of those overrings of Galois rings of integers whose class groups admit the same norm-like group action described above. This ultimately demonstrates that all class groups of interest can be realized as the class group of a simple localizations. Finally, Section 6 considers the arithmetic of norms of algebraic rings of integers—a topic closely tied to this action. We conclude by drawing an interesting connection between this arithmetic and the partition problem.
@@ 2. Direct Consequences of the Group Action
Let $p<23$ be an odd prime and $a\in\mathbb{Z}$. If $K$ is the splitting field of $x^{p}-a$, $\mathcal{O}_{K}$ is an HFD if and only if $\mathcal{O}_{K}$ is a UFD.
@@
We begin by noting that $a=1$ is a degenerative case in which the theorem holds trivially. Now, as the splitting field of $x^{p}-a$, $K$ is Galois over $\mathbb{Q}$. Furthermore, we have $K\cong\mathbb{Q}(\omega,\alpha)$ where $\omega$ is a primitive $p^{th}$ root of unity and $\alpha$ a root of $x^{p}-a$. This gives us the following lattice.
@@
Recall that the action of $G=\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$ induces a homomorphism from $G$ to $\text{Aut}(Cl_{K})$ given by $\sigma\overset{\psi}{\mapsto}\phi$ where $\phi([I])=\sigma\cdot[I]=[\sigma(I)]$. Note, the kernel of this homomorphism is precisely the kernel of the action. In this section, we will see how this map can be used to further restrict the structure of $Cl_{K}$ for a Galois number field $K$. Conversely, we will use this tool to help answer questions related to the inverse class group problem. This terminology is used in the spirit of the inverse Galois problem. In particular, the inverse class group problem asks if an arbitrary abelian group can be realized as the class group of a ring of integers. This question was famously answered in the positive for the more general class of Dedekind domains by Claborn in [9]. In the case of imaginary quadratics, the answer is known to be no. Certain groups can be shown by brute force not to appear, the smallest among them being $(\mathbb{Z}/3\mathbb{Z})^{3}$. For details, the reader is encouraged to see [25]. The question remains open for rings of integers in general. In this section, for a given finite abelian group $A$, we will explore which rings of integers may admit $Cl_{K}\cong A$. We begin with a direct proof of another result from [10] which we will return to with some new found tools in section 4.
@@
4. Localization and the Class Group
In this section, we will use localization to strengthen some of the results developed previously. First, recall that any overring—in particular, any localization—of a Dedekind domain remains Dedekind. Intuitively, as a localization of $R$ is attained by turning a set of elements $S\subseteq R$ into units, we expect the class group of $R_{S}$ to be smaller than that of $R$ as those non-principal ideals $I\subseteq R$ for which $I\cap S\neq\emptyset$ will become principal. As we will demonstrate, “small” localizations will tend to lead to small changes in the class group. In the context of rings of algebraic integers, this will be useful in further restricting the possible structure of $Cl_{K}$. The following result can be viewed as a particular case of Theorem 1.5 in [8]—one of Claborn’s pioneering works on Dedekind domains.
@@ 4. Localization and the Class Group
Let $D$ be a Dedekind domain and $\alpha\in D$ a nonzero nonunit with
$(\alpha)=\mathfrak{p}_{1}^{n_{1}}\mathfrak{p}_{2}^{n_{2}}\dotsm\mathfrak{p}_{k}^{n_{k}}.$
Then $Cl(D[\frac{1}{\alpha}])\cong Cl(D)/\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$>.
Note, $D[\frac{1}{\alpha}]$ is the localization of $D$ at the multiplicative set $\{\alpha^{n}\,|\,n\in\mathbb{N}\}$ which is the complement of the set-theoretic union of all prime ideals not containing $\alpha$.
@@
Let $\psi:Cl(D)\rightarrow Cl(D[\frac{1}{\alpha}])$ be the canonical map $I]\overset{\psi}{\mapsto}\left[ID[\frac{1}{\alpha}]\right]$. By Propositions 1.3 and 1.5 of [8], $\psi$ maps $Cl(D)$ onto $Cl(D[\frac{1}{\alpha}])$ and the kernel is generated by the classes $\mathfrak{p}_{\beta}]$ where $\mathfrak{p}_{\beta}$ ranges over the prime ideas such that $\mathfrak{p}_{\beta}D[\frac{1}{\alpha}]=D[\frac{1}{\alpha}]$. As $D[\frac{1}{\alpha}]=D_{S}$ where $S=\{\alpha^{n}\,|\,n\in\mathbb{N}\}$, these are precisely the prime ideals $\mathfrak{p}$ such that $\mathfrak{p}\cap S\neq\emptyset$ or the primes containing $\alpha^{n}$ for some $n\in\mathbb{N}$. By observing that these are the prime divisors of $(\alpha)$, we obtain the result by the first isomorphism theorem.
@@
Notably, given $Cl(D)$ is torsion, Theorem 4.1 implies that any homomorphic image of $Cl(D)$ can be realized as the class group of such a localization of $D$. Claborn gives a similar, slightly weaker result in [8] where, given a Dedekind domain $A$ with arbitrary class group $G$ and subgroup $H$, he constructs a Dedekind polynomial ring over $A$ with class group isomorphic to $G/H$. As he answered the inverse class group problem for Dedekind domains, Theorem 4.1 brings us a step closer for overrings of $\mathcal{O}_{K}$. In essence, $\mathcal{O}_{K}[\frac{1}{\alpha}]$ is very nearly a ring of integers itself. In fact, it is the precise analog if the base ring $\mathbb{Z}$ were replaced by $\mathbb{Z}[\frac{1}{N(\alpha)}]$ for some nonzero $\alpha\in\mathcal{O}_{K}$.
Theorem 4.2.
Let $K$ be a Galois number field and $\alpha\in\mathcal{O}_{K}$ a nonzero element. Then $\mathcal{O}_{K}[\frac{1}{\alpha}]$ is integral over $\mathbb{Z}[\frac{1}{N(\alpha)}]$.
Proof.
Note, it is sufficient to show $\frac{1}{\alpha}$ is integral over $\mathbb{Z}[\frac{1}{N(\alpha)}]$. Let $f(x)\in\mathbb{Z}[x]$ be the minimal polynomial of $\alpha$. Noting that $N(\alpha)$ is the product of the roots of $f(x)$, we write $f(x)=x^{n}+m_{n-1}x^{n-1}+\dotsm+m_{1}x+N(\alpha)$. Thus,
$N(\alpha)+m_{1}\alpha+\dotsm+m_{n-1}\alpha^{n-1}+\alpha^{n}=0.$
Multiplying by $\frac{1}{N(\alpha)\alpha^{n}}$ yields
$\left(\frac{1}{\alpha}\right)^{n}+\frac{m_{1}}{N(\alpha)}\left(\frac{1}{\alpha}\right)^{n-1}+\dotsm+\frac{m_{n-1}}{N(\alpha)}\left(\frac{1}{\alpha}\right)+\frac{1}{N(\alpha)}=0.$
This completes the proof. ∎
From this, we can see that $\mathcal{O}_{K}[\frac{1}{\alpha}]$ is in fact the integral closure of $\mathbb{Z}[\frac{1}{N(\alpha)}]$ in $K$ as the former is integrally closed.
$\mathbb{Q}$}K{K}[$\mathbb{Z}\!\left[\frac{1}{N(\alpha)}\right]$$\mathcal{O}_{K}\!\left[\frac{1}{\alpha}\right]$}
Note also that $Cl(\mathcal{O}_{K}[\frac{1}{\alpha}])$ is finite and retains the useful property that each class contains a prime ideal. This follows directly from the correspondence between prime ideals of $\mathcal{O}_{K}[\frac{1}{\alpha}]$ and prime ideals of $\mathcal{O}_{K}$ disjoint from $\{\alpha^{n},\,n\in\mathbb{N}\}$. This allows for the application of many theorems on rings of integers to be applied to $\mathcal{O}_{K}[\frac{1}{\alpha}]$. The following gives one such example.
@@
The remainder of this section will be dedicated to leveraging this fact to extend the techniques and theorems developed in sections 2 and 3. Our hope is that the Galois group of $K$ will act on the class group of $\mathcal{O}_{K}[\frac{1}{\alpha}]$ in the same manner as $Cl_{K}$ which would allow us to place the same restrictions on the homomorphic images of the class group—further constraining the structure of $Cl_{K}$. Unfortunately, the action $\sigma\cdot[\mathfrak{P}]=[\sigma(\mathfrak{P})]$ for $\sigma\in G$ and $\mathfrak{P}]\in Cl(\mathcal{O}_{K}[\frac{1}{\alpha}])$ will not be well-defined in general as $\sigma(\frac{1}{\alpha})=\sigma(\alpha)^{-1}$ need not be in $\mathcal{O}_{K}[\frac{1}{\alpha}]$ in general. However, in the case that $\alpha\in\mathbb{Z}$, we avoid this issue as we will have $\sigma(\frac{1}{\alpha})=\frac{1}{\alpha}$ for all $\sigma\in G$.
@@
Let $K$ be a Galois number field and $A$ the subgroup of $\text{Aut}(Cl_{K})$ induced by the action of of $\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$. Suppose that $\alpha\in\mathcal{O}_{K}$ is a nonzero nonunit and
$(\alpha)=\mathfrak{p}_{1}^{n_{1}}\mathfrak{p}_{2}^{n_{2}}\dotsm\mathfrak{p}_{k}^{n_{k}}.$
Then $Cl(\mathcal{O}_{K}[\frac{1}{N(\alpha)}])\cong Cl_{K}/A\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$> where $N(\alpha)$ is the norm of $\alpha$, and $A\left<\alpha\right$> denotes the subgroup generated by the $A$-automorphic images of the elements in $\alpha$.
@@
Recall that the ideal norm agrees with the element norm for principal ideals. Thus, as $K$ is Galois, we have $(N(\alpha))=N((\alpha))=\prod_{\sigma\in G}\sigma((\alpha))=\prod_{\sigma\in G}\sigma(\mathfrak{p}_{1})^{n_{1}}\dotsm\sigma(\mathfrak{p}_{k})^{n_{k}}$. The theorem then follows by applying Theorem 4.1 to $N(\alpha)$. ∎
Now, the group action of $G$ on the class group of $\mathcal{O}_{K}[\frac{1}{N(\alpha)}]$ will be well-defined. More than this, it will be a norm-like Galois action.
@@
For a Galois number field $K$ and nonzero rational integer $n$, the group action $\sigma\cdot[\mathfrak{P}]=[\sigma(\mathfrak{P})]$ for $\sigma\in\text{Gal}(F/\mathbb{Q}),[\mathfrak{P}]\in Cl\left(\mathcal{O}_{K}[\frac{1}{n}]\right)$ is a well-defined, norm-like action.